OpenStudy (anonymous):

what is y-5=4/5(x-5)

OpenStudy (mokeira):

do you mean something like this? \[y-5=\frac{ 4 }{5(x-5) }\]

OpenStudy (anonymous):

Is this what you mean?\[y-5=\frac{ 4 }{ 5 }(x-5)\]

OpenStudy (anonymous):

It's the equation of a line. Move -5 to the RHS and simplify.\[y=\frac{4 x}{5}+1 \]